OpenStudy (anonymous):

What is the equation of a line that passes through the point (-3,-11) and is parallel to the line who equation is 2x-y=4?

OpenStudy (anonymous):

\[2x-y=4\] \[y=2x-4\] line with slope 2, y intercept -4 if the line line you want it parallel it must have the same slope, so use the 'point-slope' formula \[y-y_1=m(x-x_1)\] so \[y=(-11)=2(x-(-3))\] \[y+11=2(x+3)\] \[y+11=2x+6\] \[y=2x-5\]

OpenStudy (anonymous):

oops typo. first line should be \[y-(-11)=2(x-(-3))\]